find an equation for the line parallel to the line 4x-3y=-9 having the same y-intercept as 7x+4y=7
Answer by blwinbbbles(106)

First convert both equations to slope intercept form which is
y = mx + b with m being the slope and b is the y intercept
Equation 1:
4x-3y=-9 subtract 4x from both sides
-3y = -4x - 9 divide both sides by -3
y = (4/3)x + 3 so the slope in equation 1 is 4/3
Equation 2:
7x + 4y = 7 subtract 7x from both sides
4y = -7x + 7 divide both sides by 4
y = (-7/4)x + 7/4 so the y-intercept is 7/4
Now parallel lines have the same slope.
the problem wants us to find the equation using the slope from equation 1 and the y intercept from equation 2
therefore the new equation will be: using the form y = mx + b
So: y = (4/3)x + 7/4